# DustNET: enabling machine learning and AI models of dusty plasmas

### 23 Mar 2026 v2

@@ IV.9 Solar and lunar dusty plasmas
Figure 27: A schematic of the circumlunar dusty plasma environment generated by ChatGPT. Dust grains are generated primarily by micrometeoroid impacts and electrostatic lofting from the lunar surface, forming a tenuous exosphere with density decreasing with altitude. The Moon is immersed in the solar wind plasma (typical density $\sim 10^{6}$ cm-3), which influences dust charging and transport. Enhanced dust activity is expected near the lunar terminator due to electric field gradients.
@@ IV.10 Synthetic data
As an example, synthetic image datasets were constructed using the DeepTrack simulation framework Midtvedt et al. (2021), which reproduces particle distributions in complex plasma environments. To improve robustness, images were generated across a range of signal-to-noise ratios (SNR), allowing the neural network to learn noise-invariant features. This strategy enhances model generalization and resilience when applied to real experimental data, where noise levels and imaging conditions can vary significantly.
@@
Beyond standard isotropic particle distributions, artificially generated string-like structures were incorporated to train the model to recognize more complex formations. Special labeled masks were created specifically for these structures to improve training effectiveness. To enhance visualization and analysis, the image resolution was increased to 256×256 pixels.
These artificial datasets played a crucial role in training a U-Net model, enabling it to detect particles and structures under varying conditions, including different noise levels and distribution complexities. By incorporating controlled variations into the synthetic data, the model was trained to generalize across a broad range of scenarios, preparing it for the challenges posed by the noisier and more diverse experimental datasets collected from the PK-4 experiment.
@@
For the study of plasma crystals, three-dimensional data was required, leading to the creation of artificial 3D datasets that accurately represent crystalline structures. These datasets modeled the spatial arrangements of hexagonal close-packed (hcp), body-centered cubic (bcc), and face-centered cubic (fcc) structures Dormagen et al. (2024b). To ensure model robustness across various experimental conditions, noise was artificially introduced at different levels, incorporating drift movements of individual particles to better simulate real experimental conditions. To introduce noise into the artificial data, Gaussian noise was added to the dataset. This Gaussian-distributed shift mimics the Brownian motion of the particles. The standard deviation of the Gaussian noise was gradually increased from to in steps of , where is the average nearest-neighbor distance. Consequently, domains were generated for each structure within a specified volume for each noise level.
@@ IV.11 Text data about dusty plasmas
Text corpora constitute an essential component of any dusty-plasma AI model, noticeably a class of models known as Foundational Models (FoMos), see Sec. (V.7) additional details. A FoMo will be enable many automation tasks, including but not limited to multi-modal data processing and interpretation, cross-referencing observational results with theoretical models, and automated literature synthesis for multi-scale phenomena. Large language models (LLMs) have emerged as a powerful FoMo framework for extracting structured knowledge, identifying trends, mapping citation networks, and linking observational datasets to governing physical principles. In the context of dusty plasmas, textual data provide not only descriptive and theoretical content but also methodological details about instrumentation, calibration, numerical modeling, and experimental protocols that are difficult to infer from numerical data alone.
The peer-reviewed dusty-plasma literature forms a well-defined domain-specific corpus. A typical scientific article has a digital size of approximately 1–5 MB (including figures, tables, and supplementary material) and it is estimated that the total number of existing dusty-plasma–relevant papers is on the order of $10^{4}$. The aggregate volume of peer-reviewed text would therefore lie in the range of 10–50 GB. In addition, monographs and edited volumes—estimated to number on the order of $\sim$50 major books in plasma physics, complex plasmas, and related topics, each potentially $\sim$100 MB in digital form—would contribute an additional $\sim$5 GB. While modest compared to general web-scale corpora, this domain-specific literature is dense in physical content and contains high-value structured information (equations, parameter tables, scaling laws, instrument descriptions, and experimental constraints) that is particularly suitable for domain adaptation of scientific LLMs.
@@ V.2 GNN model
Graph neural networks are sets of geometric deep-learning algorithms that learn from graphs. A graph represents a set of nodes (pieces of data), and the relationships between them, known as edges. GNNs were introduced in the early 2000s, physicists didn’t start using them to analyze experimental data until 2019. Unlike convolutional neural networks, which work only with data that’s presented in a regular grid, GNNs can analyze datasets that have irregular, 3D geometric shapes. A novel variant of GNN for particle tracking called Hierarchical Graph Neural Network (HGNN) was described LCF:2023. The Exa.TrkX project has applied geometric learning concepts such as metric learning and graph neural networks (GNNs) to HEP particle tracking JMC:2021. The training dataset was the TrackML dataset.
V.3 RNNs
@@ V.5 Additional Generative Models
Besides transformer, deep generative models, such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and diffusion models, have shown significant potential in generating realistic data samples. Generative models are useful when it is difficult to do experiments. Mushroom clouds associated with thermonuclear explosions are the largest man-made dusty plasmas to date. A few examples of computer-generated visualizations of a post-apocalyptic landscape in a nuclear winter with dusty plasma swirling in the sky is shown in Fig. 33.
Figure 33: Visualizations of a post-apocalyptic landscape under a nuclear winter with dusty plasmas swirling in the sky. The ground is barren and covered by dust fallouts from above. The scene, generated by DALLE 3, is eerie and haunting, capturing the desolate aftermath of a global dusty plasma.
@@
V.6 Other ML methods
There are other ML works focusing on particle tracking or image datasets. Dingel et al. (2021) presents AdaPT, an application for tracking spherical magnetic particles, which uses intensity-based localization and two linking algorithms (frame-by-frame linking and linear assignment) to connect particle positions across frames. Wollmann and Rohr (2021) introduces Deep Consensus Network, a deep neural network for object detection in microscopy images, utilizing a Feature Pyramid Network, Centroid Proposal Network, and anchor regularization within a Bayesian framework. Duarte and Vlimant (2022) uses Graph Neural Networks (GNNs) for particle tracking and reconstruction in high-energy physics, representing particle detector hits as graphs to leverage weight sharing and local connectivity, addressing many-to-one and one-to-many measurement complexities. Kalinin et al. (2021) discusses autonomous experiment methodologies in Electron and Scanning Probe Microscopy, analyzing pathways for automation using machine learning techniques inspired by recent successes in autonomous systems. Zhang and Piggott (2020) presents an unsupervised learning approach for Particle Image Velocimetry (PIV), constructing a loss function based on photometric loss, bidirectional flow consistency, and spatial smoothness to address challenges in obtaining reliable ground truth data. Wang et al. (2022) applies a Physics-Informed Neural Network (PINN) for reconstructing dense velocity fields from sparse PIV/PTV data, integrating the Navier–Stokes equations into the loss function to predict both velocity and pressure fields.
@@ V.7 AI foundation models for dusty plasmas
The dusty-plasma community will likely need several types of foundation models (FoMos), tailored to different scientific objectives. FoMos became popular around 2022, with the introduction of ChatGPT and other Large Language Models (LLMs). LLMs are AI engines for text processing, including text generation (1D sequential data). Other FoMos like DALL-E generate images (2D data). Transformer, which is at the center of state-of-the-art AI applications in industry, is the backbone of most FoMos. The difference between FoMos and general deep learning neural networks is scale: FoMos, with billions or more even trillions of tunable parameters Hudson et al. (2024); de Haan et al. (2025), are trained on datasets more massive than any available in the past, and they are also often trained without any pre-existing knowledge of what is in the data. The size of a one-trillion parameter model is at least 14 terabytes, which is more than 200 times the memory of one MI250X GPU (64 gigabytes) in supercomputers like Oak Ridge National Laboratory’s Frontier.
@@
Computing costs scale approximately linearly with parameter count: 7–8B models run on a single 24–48GB GPU, whereas 70B-class dense models typically require 2–6 $\times$ 80GB GPUs. Frontier systems are hosted on distributed high-memory clusters. Domain-adaptive pretraining for 7–70B open models may require multi-node GPU resources when token budgets exceed $10^{6}$ [Ref?].
@@
Table 2: A list of Frontier AI Foundation models for mult-modal data processing. There are over 300 known FoMos globally as of 2024 385.
Name Developer Parameters Dataset Hardware Tokens
[Billions] [NVIDIA] (#)
@@
ANI-1 SIR:2017a; SIR:2017b *** 124k DFT database GPU potential
GDB functions
@@
[…]
ChemBERTa CGR:2020 *** MoleculeNET 3220
Claude Anthropic 3220
CLIP OpenAI 3220
[…]
Cohere Cohere 52 B business text <<100 3800
DALLE OpenAI images
DBRX Databricks 3220
Dolly Databricks 3220
Fuyu-8B Adept <<1 closed
Gemini Google <<1 closed
[…]
Granite IBM 330
Jurassic AI21 Labs 7120
[…]
SciBERT BLC:2019 ‘AI2’ Semantic Scholar 3220
SciFive PAT:2021 *** biomed corpora
[…]
StarCoder Hugging face 1011 <<100 ?
T5 Google C4
Titan Amazon
@@ VI Applications of DustNET
AI foundation model may be used to predict plasma behavior, diagnose plasma states, or optimize experimental parameters.
@@ VI.3 Dusty plasma physics for AI
The state-of-the-art AI models, such as ChatGPT 5.x, have been trained primarily on large corpora of internet-scale text and images. While such datasets have enabled remarkable advances in language and reasoning tasks, the next stage of AI development is expected to rely increasingly on data grounded in the physical world. Dusty plasmas represent a particularly rich and underexplored domain in this context. The concept of a world model, advocated by Yann LeCun and others, refers to AI systems capable of learning predictive representations of real physical environments. Because dusty plasmas exhibit observable particle dynamics, strong coupling, and multi-scale interactions that can be directly measured in both laboratory and natural settings, they offer an ideal testbed for contributing to such world models.
@@
Dusty plasmas offer a unique advantage for data-driven modeling: individual dust grains are often directly observable through high-speed optical imaging. As a result, two central problems in dusty-plasma image analysis are the determination of particle positions in each frame and the automatic tracking of particles across successive frames in a time-resolved image sequence Wang et al. (2020a). Because imaging cameras typically record two-dimensional projections of inherently three-dimensional particle motion, similar challenges arise in fluid mechanics diagnostics such as particle tracking velocimetry (PTV), particle image velocimetry (PIV), and laser speckle velocimetry (LSV). These related fields have therefore developed a wide range of algorithms for particle-based data analysis that can be adapted to dusty plasma studies. Recent work shows that deep neural networks can significantly improve inference accuracy and robustness in such tasks, although these gains may come at the cost of increased computational or hardware requirements Deng et al. (2020).

### 09 Apr 2026 v3

@@ IV.9 Solar and lunar dusty plasmas
Figure 27: A schematic illustration of the circumlunar dusty plasma environment, generated using AI-assisted tools (ChatGPT). Dust grains are produced primarily through micrometeoroid impacts and electrostatic lofting from the lunar surface, forming a tenuous dust exosphere with dust density decreasing with altitude. The Moon is embedded in the solar wind plasma (typical density $\sim 10^{6}$ cm-3 at the Moon), which governs dust charging, transport, and plasma–dust interactions. Enhanced dust activity is expected near the lunar terminator, where strong electric field gradients arise from transitions between illuminated and shadowed regions. Future lunar exploration and sustained human presence may significantly perturb the natural dusty plasma environment through surface operations, resource utilization, and artificial plasma and dust generation.
@@ IV.10 Synthetic data
As an example, synthetic image datasets were constructed using the DeepTrack simulation framework Midtvedt et al. (2021), which reproduces particle distributions in complex plasma environments. To improve robustness, images were generated across a range of signal-to-noise ratios (SNR), Fig. 28, allowing the neural network to learn noise-invariant features. This strategy enhances model generalization and resilience when applied to real experimental data, where noise levels and imaging conditions can vary significantly.
@@
Beyond standard isotropic particle distributions, artificially generated string-like structures were incorporated to train the model to recognize more complex formations. Special labeled masks were created specifically for these structures to improve training effectiveness. To enhance visualization and analysis, the image resolution was increased to $\times256$ pixels.
These artificial datasets played a crucial role in training a U-Net model, enabling it to detect particles and structures under varying conditions, including different noise levels and distribution complexities, Fig. 29. By incorporating controlled variations into the synthetic data, the model was trained to generalize across a broad range of scenarios, preparing it for the challenges posed by the noisier and more diverse experimental datasets collected from the PK-4 experiment.
@@
For the study of plasma crystals, three-dimensional data was required, leading to the creation of artificial 3D datasets that accurately represent crystalline structures. These datasets modeled the spatial arrangements of hexagonal close-packed (hcp), body-centered cubic (bcc), and face-centered cubic (fcc) structures Dormagen et al. (2024b). To ensure model robustness across various experimental conditions, noise was artificially introduced at different levels, incorporating drift movements of individual particles to better simulate real experimental conditions. To introduce noise into the artificial data, Gaussian noise was added to the dataset. This Gaussian-distributed shift mimics the Brownian motion of the particles. Consequently, domains were generated for each structure within a specified volume for each noise level, Fig. 30.
@@ IV.11 Text data about dusty plasmas
Text corpora constitute an essential component of any dusty-plasma AI model, noticeably a class of models known as Foundational Models (FoMos), see Sec. (V.7) for additional details. A FoMo will enable many automation tasks, including but not limited to multi-modal data processing and interpretation, cross-referencing observational results with theoretical models, and automated literature synthesis for multi-scale phenomena. Large language models (LLMs) have emerged as a powerful FoMo framework for extracting structured knowledge, identifying trends, mapping citation networks, and linking observational datasets to governing physical principles. In the context of dusty plasmas, textual data provide not only descriptive and theoretical content but also methodological details about instrumentation, calibration, numerical modeling, and experimental protocols that are difficult to infer from numerical data alone.
The peer-reviewed dusty-plasma literature forms a well-defined domain-specific corpus. A typical scientific article has a digital size of approximately 1–5 MB (including figures, tables, and supplementary material) and it is estimated that the total number of existing dusty-plasma–relevant papers is on the order of $10^{4}$. The aggregate volume of peer-reviewed text would therefore lie in the range of 10–50 GB. In addition, monographs and edited volumes, estimated to number on the order of $\sim$50 major books in plasma physics, complex plasmas, and related topics, each potentially $\sim$100 MB in digital form, would contribute an additional $\sim$5 GB. While modest compared to general web-scale corpora, this domain-specific literature is dense in physical content and contains high-value structured information (equations, parameter tables, scaling laws, instrument descriptions, and experimental constraints) that is particularly suitable for domain adaptation of scientific LLMs.
@@ V.2 GNN model
Graph neural networks are sets of geometric deep-learning algorithms that learn from graphs. A graph represents a set of nodes (pieces of data), and the relationships between them, known as edges. GNNs were introduced in the early 2000s, physicists didn’t start using them to analyze experimental data until 2019. Unlike convolutional neural networks, which work only with data that’s presented in a regular grid, GNNs can analyze datasets that have irregular, 3D geometric shapes. A novel variant of GNN for particle tracking called Hierarchical Graph Neural Network (HGNN) was described Liu et al. (2023). The Exa.TrkX project has applied geometric learning concepts such as metric learning and graph neural networks (GNNs) to HEP particle tracking Ju et al. (2021). The training dataset was the TrackML dataset.
V.3 RNN model
@@ V.5 Additional Generative Models
Besides transformer, deep generative models, such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and diffusion models, have shown significant potential in generating realistic data samples. Generative models are useful when it is difficult to do experiments. Mushroom clouds associated with thermonuclear explosions are probably the largest man-made dusty plasmas to date. A few examples of computer-generated visualizations of a post-apocalyptic landscape in a nuclear winter with dusty plasma swirling in the sky is shown in Fig. 33.
[…]
@@
V.6 Other ML methods and regularization
Besides neural network architectures and dataset design, the formulation of loss functions plays a central role in defining machine learning (ML) methods. Incorporating physics, geometry, statistics, and scaling constraints into loss functions has led to a rich class of approaches that extend beyond purely data-driven learning. In reinforcement learning, the loss function is replaced by a reward formulation, which similarly guides model optimization and enables learning from sparse, delayed, or partially observed signals Sutton and Barto (2018); Szepesvári (2010).
Novel loss-function formulations are especially important in settings where labeled data are limited or unavailable. For example, unsupervised and self-supervised learning methods based on photometric consistency, bidirectional flow constraints, and spatial smoothness have been successfully applied to particle image velocimetry (PIV) without requiring ground-truth labels Zhang and Piggott (2020). As discussed in Sec. III.5, embedding physical knowledge directly into the learning objective further improves model fidelity and interpretability. Physics-informed approaches, for instance, reconstruct dense velocity and pressure fields from sparse PIV/PTV measurements by incorporating the Navier–Stokes equations into the loss function Wang et al. (2022). Such constraints restrict solutions to physically admissible regimes, thereby enhancing generalization, robustness, and consistency with governing laws of physics.
The diversity of loss function designs reflects the broader ML landscape, where physics-based constraints, geometric structure, and statistical regularization can be integrated in a complementary manner. These developments motivate a shift toward hybrid learning paradigms that combine data-driven flexibility with physically grounded priors. In this context, geometric and topological learning provide particularly powerful extensions for processing structured dusty plasma data. For example, AdaPT Dingel et al. (2021) provides an application for tracking spherical magnetic particles using intensity-based localization and hybrid linking strategies. Deep Consensus Networks Wollmann and Rohr (2021) employ feature pyramid architectures and Bayesian regularization for robust object detection in microscopy images. Graph-based approaches have been successfully applied in high-energy physics, where GNNs are used to reconstruct particle trajectories from detector hits, naturally handling complex measurement ambiguities Duarte and Vlimant (2022). Autonomous experimentation frameworks in electron and scanning probe microscopy Kalinin et al. (2021) demonstrate how machine learning can guide adaptive data acquisition and analysis.
Geometric and topological learning offer a complementary framework for advancing algorithms within DustNET Bronstein et al. (2021); Berzins et al. (2025), especially for particle tracking and 2D-to-3D inversion problems. Dusty plasma data are intrinsically structured: particle positions form evolving point clouds, trajectories define graphs in space–time, and collective dynamics often exhibit manifold-like organization. Conventional deep learning models, typically designed for Euclidean grids, do not fully exploit these underlying geometric and topological structures. In contrast, geometric deep learning introduces inductive biases based on symmetry, invariance, locality, and topology Glenn et al. (2025), enabling more efficient and physically meaningful representations of dust grains and complex plasma systems.
An integrated framework combining physics-informed loss functions, geometric learning, and data-driven modeling thus provides a powerful approach for analyzing dusty plasma systems. Such methods can improve tracking accuracy, enhance reconstruction fidelity, and enable cross-regime generalization, ultimately bridging experimental observations with interpretable and predictive models.
In particle tracking applications, dust grains can be naturally represented as nodes in dynamic graphs or point clouds, with edges encoding spatial proximity, interaction strength, or temporal correspondence. Graph neural networks (GNNs), Sec. V.2, and equivariant architectures Kondor (2025) leverage permutation invariance and geometric symmetries to improve robustness, particularly in dense, noisy, or partially observed environments. Topological learning methods further extend this framework by incorporating higher-order relationships, such as clusters, filaments, and crystalline ordering, through tools including simplicial complexes and persistent homology Maria et al. (2014). These approaches enable the extraction of global structural features that are not readily captured by purely local or convolutional methods.
Geometric and topological learning is also central to 2D-to-3D reconstruction, where the goal is to reconstruct three-dimensional particle trajectories from projections or time-sequence 1D or 2D dusty plasma data. By embedding geometric constraints, such as epipolar geometry Zhang (1996, 1998); Wang et al. (2016), manifold smoothness, and physical conservation laws, into neural architectures, models can learn mappings from observation space to latent 3D configurations. Neural operator frameworks and geometric deep learning methods further generalize this approach by learning continuous mappings between function spaces, enabling reconstruction of volumetric dynamics from sparse or incomplete measurements.
Within the DustNET framework, these techniques support unified representations across modalities and scales. By combining geometric and topological priors with data-driven learning, DustNET can enhance tracking accuracy, enable robust 3D reconstruction, and reveal topological signatures of collective behavior, thereby bridging experimental observations with physically interpretable models.
@@ V.7 AI foundation models for dusty plasmas
[…]
@@
Computing costs scale approximately linearly with parameter count: 7–8B models run on a single 24–48GB GPU, whereas 70B-class dense models typically require 2–6 $\times$ 80GB GPUs. Frontier systems are hosted on distributed high-memory clusters. Domain-adaptive pretraining for 7–70B open models may require multi-node GPU resources when token budgets exceed $10^{6}$ Hoffmann et al. (2022).
@@
Table 2: Examples of frontier AI foundation models for multimodal and scientific data processing. Parameter sizes and training details are approximate and may vary by version.
Name Developer Params Dataset Hardware Modalities
[B]
AlphaFold DeepMind $\sim$100–200 PDB, UniProt TPU/GPU clus. protein struct.
ANI-1 Smith et al. (2017b, a) Roitberg et al. $\sim$0.1 DFT (GDB-17) GPU mol. potentials
BERT Google 0.11–0.34 BooksCorpus, Wiki TPU text
BioBERT Lee et al. (2019) DMIS Lab $\sim$0.34 PubMed, PMC GPU biomed. text
@@
ChemBERTa Chithrananda et al. (2020) IBM $\sim$0.1 MoleculeNet, PubChem GPU chem. text
Claude Anthropic >>100 (est.) web + curated GPU clus. multimodal
CLIP OpenAI $\sim$0.4 img.-text (400M) GPU image + text
@@
DBRX Databricks 132 (MoE) curated web/code GPU clus. text
DeepSeek-V3 DeepSeek-AI (2024) DeepSeek AI 671 (37 act.) 14.8T tok., multi.+code H800/A100 clus. text + code
Dolly Databricks 12 instr.-tuning data GPU text
Fuyu-8B Adept 8 multimodal web GPU image + text
Gemini Google >>100 (est.) multimodal web TPU/GPU multimodal
GPT-4 OpenAI et al. (2024) OpenAI >>1000 (est.) web + code + RLHF GPU clus. multimodal
Granite IBM 3–20 enterprise data GPU text
Jurassic-2 AI21 Labs 7–178 web corpora GPU text
LLaMA Touvron et al. (2023) Meta 7–65 web corpus GPU text
Luminous Aleph Alpha 13–70 multi. corpora GPU text
Mistral Mistral AI 7 web + code GPU text
Mixtral Mistral AI 46 (MoE) web + code GPU clus. text
MPT-7B MosaicML 7 curated web GPU text
Palmyra Writer $\sim$5–20 business text GPU text
Phi-2 Microsoft 2.7 synthetic + curated GPU reasoning
RedPajama Together 7–65 RedPajama GPU text
SciBERT Beltagy et al. (2019) AI2 $\sim$0.11 Semantic Scholar GPU sci. text
SciFive Phan et al. (2021) Google/NIH $\sim$0.77 PubMed TPU/GPU biomed. text
Sora OpenAI undisclosed video datasets GPU clus. video
Stable Diffusion Stability AI (2022) Stability AI $\sim$0.9 LAION-5B GPU image
StarCoder HuggingFace 15.5 code datasets GPU code
T5 Google 0.22–11 C4 TPU text
Titan Amazon undisclosed enterprise/web GPU multimodal
@@ VI Applications of DustNET
Having outlined the conceptual framework, datasets, and methodological foundations of DustNET, we now turn to its potential scientific and technological applications. The DustNET ecosystem, combining curated multi-modal datasets, physics-informed machine learning, and emerging AI foundation models, provides a unifying platform for addressing long-standing challenges in dusty plasma research. In particular, DustNET enables systematic exploration across a wide range of problems, from the origin and fundamental physical properties of dust, to its interactions with plasma environments, and the emergence of collective structures across multiple spatial and temporal scales.
In this section, we organize these applications into several interconnected themes. We first examine open problems related to the origin, physical nature, and dynamics of dust, followed by dust–plasma interactions and environmental coupling. We then discuss collective dynamics, self-organization, and multiscale structure formation. Building on these foundations, we consider broader scientific and technological applications, including implications for space, astrophysical, and laboratory systems. Finally, we highlight cross-cutting opportunities enabled by data-driven discovery, the role of edge AI agents in real-time diagnostics and control, and the reciprocal impact of dusty plasma physics on the development of next-generation AI systems. Together, these directions illustrate how DustNET can serve as both a scientific tool and a conceptual bridge between plasma physics and data-centric methodologies.
@@ VI.3 Dusty plasma physics for AI
The state-of-the-art AI models, such as ChatGPT 5.x, have been trained primarily on large corpora of internet-scale text and images. While such datasets have enabled remarkable advances in language and reasoning tasks, the next stage of AI development is expected to rely increasingly on data grounded in the physical world. Dusty plasmas represent a particularly rich and underexplored domain in this context. The concept of a world model, advocated by Yann LeCun and others Ha and Schmidhuber (2018), refers to AI systems capable of learning predictive representations of real physical environments. Because dusty plasmas exhibit observable particle dynamics, strong coupling, and multi-scale interactions that can be directly measured in both laboratory and natural settings, they offer an ideal testbed for contributing to such world models.
@@
Dusty plasmas offer a unique advantage for data-driven modeling: individual dust grains are often directly observable through high-speed optical imaging. As a result, two central problems in dusty-plasma image analysis are the determination of particle positions in each frame and the automatic tracking of particles across successive frames in a time-resolved image sequence Wang et al. (2020). Because imaging cameras typically record two-dimensional projections of inherently three-dimensional particle motion, similar challenges arise in fluid mechanics diagnostics such as particle tracking velocimetry (PTV), particle image velocimetry (PIV), and laser speckle velocimetry (LSV). These related fields have therefore developed a wide range of algorithms for particle-based data analysis that can be adapted to dusty plasma studies. Recent work shows that deep neural networks can significantly improve inference accuracy and robustness in such tasks, although these gains may come at the cost of increased computational or hardware requirements Deng et al. (2020).